OpenStudy (anonymous):

prove (8^k) - (2^k) is divisible by 6 for n belonging to 0, 1, 2, 3, n,..n+1

OpenStudy (dumbcow):

\[8^{k} = 2^{3k}\] \[2^{3k} -2^{k} = 2^{k}(4^{k}-1)\]

OpenStudy (anonymous):

one of the proof is as follows (8^k) - (2^k) =(8- 2) ((8)^(k-1) +(8)^(k-1)(2)+(8)^(k-2)(2)^2 ...................(8)(2)^(k-2)+(2)^(k-1)) =6 ((8)^(k-1) +(8)^(k-1)(2)+(8)^(k-2)(2)^2 ...................(8)(2)^(k-2)+(2)^(k-1))

OpenStudy (anonymous):

1. bASE STEP n=0 1-1=0 2. kth step assume 8^k - 2^k is div by 6 3. k+1 step 8^k +1 - (2^(k+1)) = 8(8^k -2^k) + 8 * 2^k + (2^(k+1))= = 8(8^k -2^k) + 8 * 2^k + 2^k * 2= 8(8^k - 2^k) + 6 *2^k since (8^k - 2^k) divisble by 6 by assumption in kth step and 6*2^k is divisible by 6, it follows that the sum is divisible by 6

OpenStudy (anonymous):

using binomials (8^k) - (2^k) =((6+2))^k - (2^k) expand the first expression it is found that 6 will be one of the factor...